# Can weak-gravity, causality-violation arguments constrain modified gravity?

### v1

@@ I Introduction
For concreteness, consider the $2\to 2$ scattering of massive scalar particles via graviton exchanges. Using the Mandelstam variable $s$, the square of the sum of the momenta of the ingoing or outgoing particles, transplanckian gravitational scattering is defined by $\sqrt{s}\gg M_{\rm Pl}$ or $Gs\gg\hbar$. For fundamental quanta, this regime corresponds to the “ultrarelativistic limit,” and in such a high-energy limit, the so-called eikonal approximation is semi-classical. This point is evident after considering the opposite, subplanckian regime $Gs\ll\hbar$. In this case, the Schwarzschild radius scale $G\sqrt{s}$ is smaller than the Compton wavelength scale s/$\ell_{s}\sim\hbar/\sqrt{s}$ and so a quantum description is necessary. On the other hand, for $Gs\gg\hbar$, we have $G\sqrt{s}\gg\ell_{s}$ and a classical description is possible.
As we shall see in the next sections, we can obtain semiclassical observables from the eikonal limit of transplanckian scattering, such as the Shapiro time delay and the scattering angle. By Shapiro time delay, or simply time-delay for short, we refer to the difference in how long it takes for a wavepacket to propagate between two spacetime events with and without scattering by a massive source (as measured by asymptotic observers). The classroom example for wavepacket scattering is that which light from distant stars experiences as it grazes the surface of the Sun on its way to Earth, which led to the first experimental confirmation of Einstein’s theory of GR by Eddington [7]. A modern example of Shapiro time delay is that suffered by electromagnetic signals emitted by the Cassini spacecraft when it was near Jupiter, as the signals grazed the surface of the Sun on their way to Earth [8]. In GR, this time delay is always positive (i.e. it’s a time delay), as one intuitively expects based on causality.
@@
In this paper, we revisit the validity of these causality arguments for beyond-Einstein gravity theories. More specifically, we consider theories in which GR is modified by the addition of higher-derivative operators, suppressed by powers of some energy scale $M_{\lambda}\leq M_{\rm Pl}$. The theory is not valid at energies close to or larger than $M_{\lambda}$, since then, even higher-derivative operators would have had to be included. In the opposite regime, for energies much lower than $M_{\lambda}$, we expect that all observables will be merely corrected by the extra operators, where, in particular, these beyond-Einstein corrections never dominate the GR predictions. However, in transplanckian scattering, energies larger than $M_{\rm Pl}$ and hence $M_{\lambda}$ are considered, and this begs the question of whether the effective-field-theory picture is valid in such calculations. We find that the answer to this question is affirmative (i.e. the effective-field-theory picture is valid in transplackian scattering) provided we take the eikonal limit properly and respect its range of applicability.
@@
Essentially, causality constraints based on eikonal scattering can give a lower bound on $M_{\lambda}$, at most. Other lower bounds can be obtained from the lack of observational effects, for example in the gravitational waves emitted in the coalescence of black holes [35]. Upper bounds can only be obtained by observations, directly or indirectly, that positively identify a modification from GR, which has not yet occured in strong-field gravitational observations. Due to the scope of the eikonal approximation, the theoretical lower bound can never be indefinitely large. If one incorrectly sets the theoretical lower bound arbitrarily higher than any lower bound obtained from (the lack of) observations, then one would conclude that the theory is not valid and has to be modified. The question is then, how large can we assume the theoretical lower bound to be?
@@ II Eikonal approximation, phase shift, and time delay
Consider the $2\to 2$ scattering of massive scalar particles through graviton exchanges. In the subplanckian case, we consider an expansion in $Gs\ll 1$, corresponding to the Born approximation of the scattering problem, where recall that $s$ is the Mandelstam variable corresponding to the square of the sum of the incoming or outgoing 4-momenta (see e.g. [45, 43]). For the gravitational interaction, the Born amplitude is dominated by a single-graviton exchange diagram. In the transplanckian case, the expansion can be reorganized in powers of the ratio between the (square of the) momentum transferred and the (square of the) center of mass energy, i.e. $t/s\ll 1$ [46]. This corresponds to the eikonal approximation [47], and it is equivalent to an expansion in $G\sqrt{s}/b\ll 1$ where $b$ is the impact parameter of the scattering event. In the transplanckian case, the eikonal amplitude is dominated by graviton ladder loop diagrams, corresponding to multigraviton exchanges (see Fig. 1) [48, 49].
@@
×$\times\int d^{D-2}b_{\perp}e^{-iq_{\perp}\cdot b_{\perp}/\hbar}\left(e^{2i\delta(b_{\perp})/\hbar}-1\right),$ (1)
@@
×$\times\int\frac{d^{D-2}q_{\perp}}{(2\pi)^{D-2}}e^{iq_{\perp}\cdot b_{\perp}/\hbar}i\mathcal{M}_{0}(t=-q_{\perp}^{2}),$ (2)
@@
×$\times\frac{\Gamma\left(\frac{D-4}{2}\right)}{(\pi b^{2})^{\frac{D-4}{2}}}.$ (4)
Figure 1: The eikonal amplitude is obtained by summing multiloop, graviton-tree diagrams. The wiggly lines represent gravitons, while the double and single lines represent quanta minimally coupled to gravity. The dots represent not only higher-loop-order diagrams but also permutations of these. Only in the Born limit, can the amplitude be approximated by the tree-level, single-graviton exchange diagram.
@@
where the dots stand of terms that do not depend on $b$ and terms that vanish when $\epsilon\to 0$. These terms and the divergence do not appear in any physical observable and we shall drop them in the following. In the ultrarelativistic limit (or massless) case, we have
@@
The second term represents the outgoing wavepacket, so performing a similar caculation as above, we have (for $t\to\infty$),
@@ III Causality constraints from gravitational scattering
which coincides with the leading-order contribution (in a weak-field expansion) to the Shapiro time delay computed by a null geodesic traveling past a Schwarzschild black hole of mass $M$, if we set $2\sqrt{s}\to M$ (see discussion in Sec. V.1). In the ultrarelativistic limit, we can also find the time delay from the geometry by means of schockwave metrics [6, 52, 53], and the resulting time delay matches the one computed from the phase shift.
The seminal paper [13] started the recent usage of time delays computed from amplitudes in modified gravity theories to constrain the coefficient of higher-order operators in the action. The idea is that extensions to GR will introduce new vertices into the S-matrix calculation. These new vertices can lead to beyond-GR corrections to the time delay that can be negative and are proportional to the coupling constants. If the coupling constants are chosen to be large enough, these corrections can dominate over the GR contribution to the time delay, and thus, give rise to overall negative time delays (i.e. time advances). Certainly, this ought to not be allowed on physical grounds, which then allows one to place constraints on modified gravity from purely theoretical causality arguments. In recent years, an explosion of work has been carried out in this direction [14, 16, 17, 18, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
@@
The matter action supports a non-trivial background metric $\bar{g}_{\mu\nu}$. Now consider the traverse-traceless (TT) tensor perturbations of this metric such that $g_{\mu\nu}=\bar{g}_{\mu\nu}+h_{\mu\nu}$ where ¯$\overline{\nabla}^{\mu}h_{\mu\nu}=0=\bar{g}^{\mu\nu}h_{\mu\nu}$. The action for the metric perturbation is, up to quadratic order in $h_{\mu\nu}$,
@@
+¯$+\left.\overline{\nabla}_{\alpha}\overline{\nabla}_{\beta}h_{\nu\mu}-\overline{\nabla}_{\nu}\overline{\nabla}_{\beta}h_{\alpha\mu}+\overline{\nabla}_{\nu}\overline{\nabla}_{\mu}h_{\alpha\beta}\right)h^{\mu\nu}h^{\alpha\beta}.$ (15)
We see that there are not only derivative terms but also cubic terms proportional to the background curvature tensor in the third-order perturbative expansion of the Einstein-Hilbert action. In regions where the curvature is small, we can safely neglect the terms proportional to the background curvature tensor. Similarly, since $h_{\mu\nu}$ varies significantly more than $\bar{g}$ over distances smaller than the radius of curvature, we can set ¯$\overline{\nabla}h\to\partial h$ in the weak-gravity region. For instance, in the Schwarzschild spacetime, we have the Kretschmann scalar
@@
To go beyond linearized gravity, we need to start considering graviton three- and four-vertices. Indeed, for the Schwarzchild case, the nonlinear terms can be resummed to get the full solution from amplitudes [55, 56]. For the scattering of two massive bodies, the leading-loop corrections as $b\to G\sqrt{s}$ come from graviton trees between the scalar external lines [49]. As an example, the order of magnitude of the contribution of the diagram in Fig. 3 to $\delta(b_{\perp})$ can be estimated as follows: each graviton vertex attached to an external lines contributes a factor $\kappa\sqrt{s}$ coming from the vertex ($\kappa s$) times the internal propagator ($1/\sqrt{s}$) contributions (see e.g. the estimates in [45, 44]). The vertices connecting internal lines do not have the extra $\sqrt{s}$ factor, but come with the loop and transferred momentum factors. These will give rise to the $b_{\perp}$-dependent factors in $\delta(b_{\perp})$ that can be estimated by dimensional analysis. For the diagram in Fig. 3, we have a contribution of ${\cal{O}}[\kappa^{6}(\sqrt{s})^{4}b^{-2(D-3)}]={\cal{O}}[Gs(r^{2}_{s}/b^{2})^{D-3}]$, which differs from the order of the tree-level diagram in the first line of Fig. 1 by a factor of ${\cal{O}}[(r^{2}_{s}/b^{2})^{D-3}]$.
@@
The limitation described in the previous paragraph is also discussed in [13], although in different words. In that work, the authors considered a $D>4$ modification to GR by adding the Gauss-Bonet term to the Einstein-Hilbert Lagrangian. The new terms have a coefficient $\lambda$ with dimension of length squared. Then, the leading-order correction to the GR prediction for $\delta(b_{\perp})$ for the scattering of gravitons [^2] We can apply the discussion of the diagrams of the $2\to 2$ scattering of the scalars to the scattering of gravitons after taking the scalar masses to zero and focusing on individual graviton helicities. This is the case for any theory in which the kinetic terms for these helicities are diagonal and canonically normalized, such as GR. In dCS gravity, we need to work with circular polarizations, $h_{+}\pm ih_{\times}$, for that to be the case. is suppressed by a factor that scales as $\lambda/b^{2}$ relative to the GR contribution, which is $\mathcal{O}\left({Gs}/{b^{D-4}}\right)$ [57, 58]. Moreover, this correction can overwhelm the GR prediction for the time delay, leading to the possibility of a time advance, when $b^{2}\lesssim\lambda$. We then see that the exact constraint on $\lambda$ always depends on how small a $b$ we can trust in our calculations. Although similar arguments have been discussed in [18], we are unaware of estimates on how small $b$ can be before causality arguments cease to be self-consistent. More importantly, many subsequent studies do not even mention this intrinsic limitation. In the following section, we will find a bound on the coefficients of high-order operators for the eikonal approximation to be self-consistent in the weak-gravity limit.
@@ IV.1 Regime of validity of weak-field analysis of modified gravity
We start by estimating how small the impact parameter can be before the eikonal approximation breaks down in a generic beyond-Eistein effective field theory. Consider adding a higher-order curvature operator to the Einstein-Hilbert Lagrangian, such that a new cubic graviton vertex is introduced. Suppose that it gives rise to a correction to the eikonal phase of the form
@@
Figure 4: Possible hierarchy of scales. In cases I and II, the correction to the phase shift does not introduce causality issues, since $b>\lambda_{n}^{1/n}$. For case III, we can have a time advance, but since $b\gg r_{s}$ for consistency, we should also have $\lambda_{n}^{1/n}\gg r_{s}$. This is only possible when the corrections to the Einstein-Hilbert action are understood as defining an exact theory. If the corrected action is to be understood as an effective field theory, then $\lambda_{n}^{1/n}\ll r_{s}$ for the Einstein-Hilbert action to dominate over the higher-derivative operators, which then disallows case III completely. This figure is not to scale.
@@ IV.2 Causality constraints in dynamical Chern-Simons gravity
for the +++++ + polarization, corresponding to $+2$ graviton helicity, where $q_{i}$ are the components of the transverse exchange momentum. Here $m_{\sigma}$ is the mass of the heavy field and $\omega$ is the graviton energy in the center-of-mass frame. With a --- - polarization in the external state we would get $\mathcal{M}^{*}$ for the scattering amplitude. This result implies that, if we send a graviton to probe the heavy field $\sigma$, there is a non-vanishing probability of this graviton turning into a dCS scalar quanta by exchanging a graviton.
@@
Let us then investigate how large these loop corrections to the eikonal scattering amplitudes can be. In the GR case, loop corrections to the eikonal amplitude of graviton-mediated scalar scattering have been studied in [59]. For dCS gravity, the leading diagrams are shown in Fig. 6. The diagram (A) in Fig. 6 has a dCS vertex, proportional to $\kappa\alpha$ and, by a similar power-couting argument as in the GR case, this diagram gives a contribution of

### v2

@@ I Introduction
For concreteness, consider the $2\to 2$ scattering of massive scalar particles via graviton exchanges. Using the Mandelstam variable $s$, the square of the sum of the momenta of the ingoing or outgoing particles, transplanckian gravitational scattering is defined by $\sqrt{s}\gg M_{\rm Pl}$ or $Gs\gg\hbar$. For fundamental quanta, this regime corresponds to the “ultra-relativistic limit,” and in such a high-energy limit, the so-called eikonal approximation is semi-classical. This point is evident after considering the opposite, subplanckian regime $Gs\ll\hbar$. In this case, the Schwarzschild radius scale $G\sqrt{s}$ is smaller than the Compton wavelength scale $\ell_{s}\sim\hbar/\sqrt{s}$ and so a quantum description is necessary. On the other hand, for $Gs\gg\hbar$, we have $G\sqrt{s}\gg\ell_{s}$ and a classical description is possible.
As we shall see in the next sections, we can obtain semiclassical observables from the eikonal limit of transplanckian scattering, such as the Shapiro time delay and the scattering angle. By Shapiro time delay, or simply time-delay for short, we refer to the difference in how long it takes for a wave packet to propagate between two spacetime events with and without scattering by a massive source (as measured by asymptotic observers). The classroom example for wave-packet scattering is that which light from distant stars experiences as it grazes the surface of the Sun on its way to Earth, which led to the first experimental confirmation of Einstein’s theory of GR by Eddington [7]. A modern example of Shapiro time delay is that suffered by electromagnetic signals emitted by the Cassini spacecraft when it was near Jupiter, as the signals grazed the surface of the Sun on their way to Earth [8]. In GR, this time delay is always positive (i.e. it’s a time delay), as one intuitively expects based on causality.
@@
In this paper, we revisit the validity of these causality arguments for beyond-Einstein gravity theories. More specifically, we consider theories in which GR is modified by the addition of higher-derivative operators, suppressed by powers of some energy scale $M_{\lambda}\leq M_{\rm Pl}$. The theory is not valid at energies close to or larger than $M_{\lambda}$, since then, even higher-derivative operators would have had to be included. In the opposite regime, for energies much lower than $M_{\lambda}$, we expect that all observables will be merely corrected by the extra operators, where, in particular, these beyond-Einstein corrections never dominate the GR predictions. However, in transplanckian scattering, energies larger than $M_{\rm Pl}$ and hence $M_{\lambda}$ are considered, and this begs the question of whether the effective-field-theory picture is valid in such calculations. We find that the answer to this question is affirmative (i.e. the effective-field-theory picture is valid in transplanckian scattering) provided we take the eikonal limit properly and respect its range of applicability.
@@
Essentially, causality constraints based on eikonal scattering can give a lower bound on $M_{\lambda}$, at most. Other lower bounds can be obtained from the lack of observational effects, for example in the gravitational waves emitted in the coalescence of black holes [35]. Upper bounds can only be obtained by observations, directly or indirectly, that positively identify a modification from GR, which has not yet occurred in strong-field gravitational observations. Due to the scope of the eikonal approximation, the theoretical lower bound can never be indefinitely large. If one incorrectly sets the theoretical lower bound arbitrarily higher than any lower bound obtained from (the lack of) observations, then one would conclude that the theory is not valid and has to be modified. The question is then, how large can we assume the theoretical lower bound to be?
@@ II Eikonal approximation, phase shift, and time delay
Consider the $2\to 2$ scattering of massive scalar particles through graviton exchanges. In the subplanckian case, we consider an expansion in $Gs\ll 1$, corresponding to the Born approximation of the scattering problem, where recall that $s$ is the Mandelstam variable corresponding to the square of the sum of the incoming or outgoing 4-momenta (see e.g. [45, 43]). For the gravitational interaction, the Born amplitude is dominated by a single-graviton exchange diagram. In the transplanckian case, the expansion can be reorganized in powers of the ratio between the (square of the) momentum transferred and the (square of the) center of mass energy, i.e. $t/s\ll 1$ [46]. This corresponds to the eikonal approximation [47], and it is equivalent to an expansion in $G\sqrt{s}/b\ll 1$ where $b$ is the impact parameter of the scattering event. In the transplanckian case, the eikonal amplitude is dominated by graviton ladder loop diagrams, corresponding to multi-graviton exchanges (see Fig. 1) [48, 49].
@@
$\times\int d^{D-2}b_{\perp}e^{-iq_{\perp}\cdot b_{\perp}/\hbar}\left(e^{2i\delta(b_{\perp})/\hbar}-1\right),$ (1)
@@
$\times\int\frac{d^{D-2}q_{\perp}}{(2\pi)^{D-2}}e^{iq_{\perp}\cdot b_{\perp}/\hbar}i\mathcal{M}_{0}(t=-q_{\perp}^{2}),$ (2)
@@
$\times\frac{\Gamma\left(\frac{D-4}{2}\right)}{(\pi b^{2})^{\frac{D-4}{2}}}.$ (4)
Figure 1: The eikonal amplitude is obtained by summing multi-loop, graviton-tree diagrams. The wiggly lines represent gravitons, while the double and single lines represent quanta minimally coupled to gravity. The dots represent not only higher-loop-order diagrams but also permutations of these. Only in the Born limit, can the amplitude be approximated by the tree-level, single-graviton exchange diagram.
@@
where the dots stand of terms that do not depend on $b$ and terms that vanish when $\epsilon\to 0$. These terms and the divergence do not appear in any physical observable and we shall drop them in the following. In the ultra-relativistic limit (or massless) case, we have
@@
The second term represents the outgoing wave packet, so performing a similar calculation as above, we have (for $t\to\infty$),
@@ III Causality constraints from gravitational scattering
which coincides with the leading-order contribution (in a weak-field expansion) to the Shapiro time delay computed by a null geodesic traveling past a Schwarzschild black hole of mass $M$, if we set $2\sqrt{s}\to M$ (see discussion in Sec. V.1). In the ultra-relativistic limit, we can also find the time delay from the geometry by means of shockwave metrics [6, 52, 53], and the resulting time delay matches the one computed from the phase shift.
The seminal paper [13] started the recent usage of time delays computed from amplitudes in modified gravity theories to constrain the coefficient of higher-order operators in the action. The idea is that extensions to GR will introduce new vertices into the S-matrix calculation. These new vertices can lead to beyond-GR corrections to the time delay that can be negative and are proportional to the coupling constants. If the coupling constants are chosen to be large enough, these corrections can dominate over the GR contribution to the time delay, and thus, give rise to overall negative time delays (i.e. time advances). Certainly, this ought to not be allowed on physical grounds, which then allows one to place constraints on modified gravity from purely theoretical causality arguments. In recent years, an explosion of work has been carried out in this direction [^1] Some of them focus on the subplanckian regime and hence do not assume the eikonal approximation, while others include some non-perturbative GR effects (see e.g. [34, 24]). [14, 16, 17, 18, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
@@
The matter action supports a non-trivial background metric $\bar{g}_{\mu\nu}$. Now consider the traverse-traceless (TT) tensor perturbations of this metric such that $g_{\mu\nu}=\bar{g}_{\mu\nu}+h_{\mu\nu}$ where $\overline{\nabla}^{\mu}h_{\mu\nu}=0=\bar{g}^{\mu\nu}h_{\mu\nu}$. The action for the metric perturbation is, up to quadratic order in $h_{\mu\nu}$,
@@
$+\left.\overline{\nabla}_{\alpha}\overline{\nabla}_{\beta}h_{\nu\mu}-\overline{\nabla}_{\nu}\overline{\nabla}_{\beta}h_{\alpha\mu}+\overline{\nabla}_{\nu}\overline{\nabla}_{\mu}h_{\alpha\beta}\right)h^{\mu\nu}h^{\alpha\beta}.$ (15)
We see that there are not only derivative terms but also cubic terms proportional to the background curvature tensor in the third-order perturbative expansion of the Einstein-Hilbert action. In regions where the curvature is small, we can safely neglect the terms proportional to the background curvature tensor. Similarly, since $h_{\mu\nu}$ varies significantly more than $\bar{g}$ over distances smaller than the radius of curvature, we can set $\overline{\nabla}h\to\partial h$ in the weak-gravity region. For instance, in the Schwarzschild spacetime, we have the Kretschmann scalar
@@
To go beyond linearized gravity, we need to start considering graviton three- and four-vertices. Indeed, for the Schwarzschild case, the nonlinear terms can be resummed to get the full solution from amplitudes [55, 56]. For the scattering of two massive bodies, the leading-loop corrections as $b\to G\sqrt{s}$ come from graviton trees between the scalar external lines [49]. As an example, the order of magnitude of the contribution of the diagram in Fig. 3 to $\delta(b_{\perp})$ can be estimated as follows: each graviton vertex attached to an external lines contributes a factor $\kappa\sqrt{s}$ coming from the vertex ($\kappa s$) times the internal propagator ($1/\sqrt{s}$) contributions (see e.g. the estimates in [45, 44]). The vertices connecting internal lines do not have the extra $\sqrt{s}$ factor, but come with the loop and transferred momentum factors. These will give rise to the $b_{\perp}$-dependent factors in $\delta(b_{\perp})$ that can be estimated by dimensional analysis. For the diagram in Fig. 3, we have a contribution of ${\cal{O}}[\kappa^{6}(\sqrt{s})^{4}b^{-2(D-3)}]={\cal{O}}[Gs(r^{2}_{s}/b^{2})^{D-3}]$, which differs from the order of the tree-level diagram in the first line of Fig. 1 by a factor of ${\cal{O}}[(r^{2}_{s}/b^{2})^{D-3}]$.
@@
The limitation described in the previous paragraph is also discussed in [13], although in different words. In that work, the authors considered a $D>4$ modification to GR by adding the Gauss-Bonnet term to the Einstein-Hilbert Lagrangian. The new terms have a coefficient $\lambda$ with dimension of length squared. Then, the leading-order correction to the GR prediction for $\delta(b_{\perp})$ for the scattering of gravitons [^3] We can apply the discussion of the diagrams of the $2\to 2$ scattering of the scalars to the scattering of gravitons after taking the scalar masses to zero and focusing on individual graviton helicities. This is the case for any theory in which the kinetic terms for these helicities are diagonal and canonically normalized, such as GR. In dCS gravity, we need to work with circular polarizations, $h_{+}\pm ih_{\times}$, for that to be the case. is suppressed by a factor that scales as $\lambda/b^{2}$ relative to the GR contribution, which is $\mathcal{O}\left({Gs}/{b^{D-4}}\right)$ [57, 58]. Moreover, this correction can overwhelm the GR prediction for the time delay, leading to the possibility of a time advance, when $b^{2}\lesssim\lambda$. We then see that the exact constraint on $\lambda$ always depends on how small a $b$ we can trust in our calculations. Although similar arguments have been discussed in [18], we are unaware of estimates on how small $b$ can be before causality arguments cease to be self-consistent. More importantly, many subsequent studies do not even mention this intrinsic limitation. In the following section, we will find a bound on the coefficients of high-order operators for the eikonal approximation to be self-consistent in the weak-gravity limit.
@@ IV.1 Regime of validity of weak-field analysis of modified gravity
We start by estimating how small the impact parameter can be before the eikonal approximation breaks down in a generic beyond-Einstein effective field theory. Consider adding a higher-order curvature operator to the Einstein-Hilbert Lagrangian, such that a new cubic graviton vertex is introduced. Suppose that it gives rise to a correction to the eikonal phase of the form
@@
Figure 4: Possible hierarchy of scales. In cases I and $I$, the correction to the phase shift does not introduce causality issues, since $b>\lambda_{n}^{1/n}$. For case III, we can have a time advance, but since $b\gg r_{s}$ for consistency, we should also have $\lambda_{n}^{1/n}\gg r_{s}$. This is only possible when the corrections to the Einstein-Hilbert action are understood as defining an exact theory. If the corrected action is to be understood as an effective field theory, then $\lambda_{n}^{1/n}\ll r_{s}$ for the Einstein-Hilbert action to dominate over the higher-derivative operators, which then disallows case III completely. This figure is not to scale.
@@ IV.2 Causality constraints in dynamical Chern-Simons gravity
for the ++++ polarization, corresponding to +2+2 graviton helicity, where $q_{i}$ are the components of the transverse exchange momentum. Here $m_{\sigma}$ is the mass of the heavy field and $\omega$ is the graviton energy in the center-of-mass frame. With a -- polarization in the external state we would get $\mathcal{M}^{*}$ for the scattering amplitude. This result implies that, if we send a graviton to probe the heavy field $\sigma$, there is a non-vanishing probability of this graviton turning into a dCS scalar quanta by exchanging a graviton.
@@
Let us then investigate how large these loop corrections to the eikonal scattering amplitudes can be. In the GR case, loop corrections to the eikonal amplitude of graviton-mediated scalar scattering have been studied in [59]. For dCS gravity, the leading diagrams are shown in Fig. 6. The diagram (A) in Fig. 6 has a dCS vertex, proportional to $\kappa\alpha$ and, by a similar power-counting argument as in the GR case, this diagram gives a contribution of